# Kolmogorov-Arnold Attention: Is Learnable Attention Better For Vision Transformers?

### v2

@@ 4 Our Architecture
Different from learning activations for MLP, as in KAN and KAT, implementing the generic learnable attention incurs impractical computational and memory costs on present computing hardware and DL toolkits. For learnable activation, $\tilde{\sigma}:\mathbb{R}^{N}\to\mathbb{R}^{N}$, a full-rank operator, ${\Phi^{i,j}\in\mathbb{R}^{N\times N}}$ acts on each row ${\cal A}^{i,j}_{k,:}$ of ${\cal A}^{i,j}$, and produces $\tilde{\sigma}({\cal A}^{i,j}_{k,:}).$ These operations are extremely compute-heavy and have large memory footprints. E.g., ViT-Tiny [24, 76, 68] has 5.53M parameters and requires nearly 0.9 GB of GPU memory to train. Implementing the learnable attention (2) in ViT-Tiny training with B-splines of order 5 and grid size 10 and Fourier basis with grid size 10 increases the parameter count to 30.68M and 39.06M, respectively. We could not train the version with B-splines for its humongous memory requirements, and B-spline computations are non-parallelizable. The one with the Fourier basis is parallelizable, but it takes approximately 60 GB of GPU memory when computing with a batch of one image. This computational bottleneck is agnostic of the basis functions.
@@
Figure 2: Different configurations to update $\widehat{\Phi}$:(a) Blockwise configuration, where $\Phi^{i,1}\neq\Phi^{i,2}\neq\cdots\neq\Phi^{i,L}$ for all $i=1,2,...,h$; (b) universal configuration, where $\Phi^{i,1}=\Phi^{i,2}=\cdots=\Phi^{i,L}=\Phi^{i}$ for all $i=1,2,...,h.$
@@
How can we remedy this? Deep neural networks exhibit low-rank structure [60, 36]. Recently, [41] showed that across various networks, the training largely occurs within a low-dimensional subspace. Therefore, we postulate that the attention matrices also have an underlying low-rank structure.
@@
To validate this, we perform spectral analysis of the attention matrices before and after the softmax activation; see Figures 6-6. The scree test [16] shows that for traditional attention matrices without softmax, there are 8 significant singular values; after softmax activation, this count increases to 16. The above observation verifies that with or without the nonlinear activation, the attention matrix $\mathcal{A}^{i,j}$ possesses an underlying low-rank structure. This motivates us to use a lower-dimensional operator, ${\widehat{\Phi}}$ for the learnable attention calculation with $r=12$. In §B.3.2-B.3.3, we perform an ablation to find the best $r.$
@@
Instead of ${\Phi^{i,j}\in\mathbb{R}^{N\times N}}$, we use an operator, ${\widehat{\Phi}^{i,j}\in\mathbb{R}^{r\times N}}$ such that ${r\ll N}$, and the new learned activation is $r$-dimensional for $k\in[N]$. This process significantly reduces the computational overhead. Next, we post-multiply another learnable weight matrix, $W^{i,j}\in\mathbb{R}^{N\times r}$ to project them back to their original dimension. For each ${k\in[N]}$, this operation results in computing ${\widehat{\sigma}({\cal A}^{i,j}_{k,:})=\left[W^{i,j}\widehat{\Phi}^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\right]^{\top}}$ (Figure 1(b)(iv)); see other variants in §B.4. Our approximation is not unique or the best; it is one of the feasible solutions given the present computing interface.
@@
Fourier Kolmogorov-Arnold Attention. Although the default basis for KANs is B-Splines in [51], the number of MHSA parameters is much more than that of an MLP. Specifically, for $L$ encoder blocks, each with $h$ attention heads, the parameter complexity is $O(N^{2}Lh(G+k))$, for $\Phi\in\mathbb{R}^{N\times N}$ and $O(NrLh(G+k))$, for $\widehat{\Phi}\in\mathbb{R}^{r\times N}$, if the model uses B-Splines of degree $k$.
@@
Moreover, as [80] mentioned, B-splines are localized functions, not standard CUDA functions. Although efficient CUDA implementations for cubic B-Splines exist [66, 65], their overall implementation results in slower, sparse, non-scalable, and complicated GPU execution; also, see [78, 61]. In §B.3.1, our extensive experiments with B-spline basis in KAN for image classification tasks verify their poor generalizability on medium-scale datasets (e.g., CIFAR-10 and CIFAR-100).
@@
[14] Let $f$ be an ${L^{p}}$ periodic function for ${p\in(1,\infty]}$, with Fourier coefficients ${\hat{f}(n)}$. Then ${\lim_{N\to\infty}\sum_{|n|\leq N}{\hat{f}}(n)e^{inx}=f(x)},$ for almost every $x$.
In particular, if a function is continuously differentiable, its Fourier series converges to it everywhere. In the past, [78, 53] used the Fourier basis in KAN, [22] proposed a Fourier analysis network (FAN) to capture periodic phenomena. Motivated by these, we use the Fourier basis to approximate the effect of the smooth softmax function. We employ the Fourier basis, $\{\sin{(\cdot)},\cos{(\cdot)}\}$ with gridsize $G$ to design learnable attention. Algorithm 1 can be used optionally. For the operator ${\widehat{\Phi}^{i,j}=[\widehat{\phi}^{i,j}_{pq}]\in\mathbb{R}^{r\times N}}$, we have ${\widehat{\Phi}^{i,j}:\mathbb{R}^{N}\to\mathbb{R}^{r}}$. Hence, each row ${{\cal A}^{i,j}_{k,:}}$, for ${k\in[N]}$, transformed into ${\widehat{\Phi}^{i,j}_{p}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]=\sum_{q=1}^{N}\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})},$ via the Fourier bases such that
@@
where $m$ refers to the ${m^{\rm th}}$ grid point on a uniform grid of size $G$. The weights, [{${[\{[a_{pqm}],[b_{pqm}]\}_{m=1}^{G},W^{i,j}]}$ are updated in the backpropagation. If we use a Fourier basis in ${\widehat{\Phi}\in\mathbb{R}^{r\times N}}$, the parameter complexity of MHSA becomes ${O(2NrhGL)}.$ Regardless of the basis used, the linear projection takes $O(N^{2}rhL)$ FLOPs. For any Transformer, $L$ and $h$ are fixed. So, one may omit them from the complexity results. See FLOPs and memory requirement in Table 12.
@@
+ $G_{1}B$ 81.81 (1.97$\downarrow$) 99.01 55.92(3.70$\downarrow$) 82.04 68.03(6.68$\downarrow$) 86.41 87.51M (1.98% $\uparrow$)
+ $G_{1}U$ 80.75(3.24$\downarrow$) 98.76 57.36 (1.22% $\downarrow$) 82.89 68.83 (5.58$\downarrow$) 87.69 85.95M (0.16% $\uparrow$)
@@
+ $G_{3}B$ 79.78 (1.60$\downarrow$) 98.70 54.11 (1.20$\uparrow$) 81.02 67.77 (3.87$\downarrow$) 87.51 23.58M (6.94$\uparrow$)
+ $G_{3}U$ 79.52(1.92$\downarrow$) 98.85 53.86(0.73$\uparrow$) 81.45 67.76(3.89$\downarrow$) 87.60 22.18M (0.56$\uparrow$)
@@
+ $G_{3}B$ 76.69(5.40$\uparrow$) 98.57 46.29(6.34$\uparrow$) 77.02 59.11 (0.07$\downarrow$) 82.01 6.29M (13.74$\uparrow$)
+ $G_{3}U$ 75.56 (3.85$\uparrow$) 98.48 46.75(7.40$\uparrow$) 76.81 57.97(1.99$\downarrow$) 81.03 5.59M (1.08$\uparrow$)
@@
Blockwise and Universal operator configuration. We consider two configurations for updating the operator $\widehat{\Phi}^{i,j}$. (a) Blockwise: In this configuration, each encoder block learns the attention through $h$ distinct operators $\widehat{\Phi}$ for each of the $h$ heads, totaling $hL$ operators; see Figure 2 (a). Like the MHSA in vanilla ViTs, the blockwise configuration is designed to learn as many different data representations as possible. (b) Universal: This is motivated from the KAT [80]. In KAT, the MLP head is replaced with different variations of a KAN head—KAN and Group-KAN. In Group-KAN, the KAN layer shares rational base functions and their coefficients among the edges. Inspired by this, in our update configuration, all $L$ encoder blocks share the same $h$ operators; $\widehat{\Phi}^{i,j}=\widehat{\Phi}^{i}$ for $j=1,2,\ldots,L$; see Figure 2-(b). Rather than learning attention through $hL$ operators, this configuration only uses $h$ operators. Here, we share all learnable units and their parameters in each head across $L$ blocks. We postulate that the blockwise mode with more operators captures more nuances from the data. In contrast, the universal mode is suitable for learning simpler decision boundaries from the data. We also note that these two modes can be used with (2) as shown in Figure 1(b). Finally, Algorithm 2 in the Appendix gives pseudocode of applying Fourier-KArAt in the ViT encoder block. We provide more nuanced discussion about these two configurations in §B.3.6.
@@ 5 Empirical Study
Implementing KArAt. We provide the implementation details in §B.1; see §B.1.1 for baselines and datasets. We integrate 5 wavelet bases, rational basis, Fourier basis, and 3 different base activation functions in our learnable MHSA design; see §B.1.3. Fourier KArAt performs the best, and we report it in the main paper. It has 1 hyperparameter, the grid size, $G$, and 2 configurations, blockwise or universal. We vary grid sizes from the set ${G=\{1,3,5,7,9,11\}.}$ For each value, we perform universal and blockwise updates; see Figure 2. With this formalization, $G_{n}B$ and $G_{n}U$ denote Fourier-KArAt with grid size $n$ and weights updated in blockwise and universal mode, respectively. See the ablation study of these variants in §B.4. We dispense the $\ell_{1}$ projection, as projecting the learned attention vectors to a probability simplex degrades the performance; see §B.3.5.
Table 3: Performance comparison of Fourier-KArAt and softmax attention on various ViT architectures.
@@
Table 4: Fourier KArAt on object detection and instance segmentation tasks on MS COCO [48] dataset. The header Box and Mask refer to detection and segmentation tasks, respectively.
Model (Initialization) Box Mask
AP AP50 AP75 AP AP50 AP75
ViT-Det-Base (ImageNet-1K) 32.34 49.16 34.43 28.35 46.04 29.48
+ $G_{1}B$ 22.48 36.82 23.13 20.11 34.09 20.46
+ $G_{1}U$ 26.68 43.25 27.88 24.01 40.21 24.66
ViT-Det-Base (Random) 15.28 26.88 15.22 13.39 24.34 12.86
+ $G_{1}B$ 10.32 18.77 10.05 8.94 16.51 8.50
+ $G_{1}U$ 10.99 20.04 10.82 9.69 17.99 9.25
(a) ViT-Tiny
(b) ViT-Tiny+Fourier KArAt
(c) ViT-Base
(d) ViT-Base+Fourier KArAt
5.1 Training Performance, Model Quality, Transferability, and Generalizability
(i) Image Classification. Table 2 presents the performance of the best-performing Fourier KArAt variants; see Figure 7 for the training loss and test accuracy curves. Fourier KArAt is easily optimized in the initial training phase, gaining at par or better loss value and accuracy than the vanilla ViTs. However, in the later training phase, except for the fewer parameter variants (ViT-Tiny+Fourier KArAt), the loss curve flattens faster; we infer that for most of the Fourier-KArAt scenarios. We also notice that larger models like ViT-Base are more susceptible to slight changes in the parameters, making the training process challenging to manage, which we try to investigate in the rest of the paper using several analytical studies. Overall, ViT-Tiny+Fourier KArAt variants outperform ViT-Tiny on CIFAR-10 and CIFAR-100 by $5.40\%$ and $7.40\%$, respectively, and on ImageNet-1K, it achieves a similar accuracy. ViT-Small and -Base models with Fourier KArAt variants can barely outperform the vanilla models, and the accuracy differences increase as we tackle larger base models.
(ii) Performance on other ViTs for Image Classification. We fused KArAt with modern ViT-based architectures, ConViT [20] and Swin Transformer [49]; §B.1.4 provides their implementation details. Table 4 shows that Fourier KArAt outperforms ConViT on CIFAR-10, experiences a modest drop on ImageNet-1K, and shows lower performance for hierarchical models, such as Swin, which requires a careful investigation.
(iii) Transferability: Image Classification. Unlike ViTs, which were pre-trained on large datasets and then transferred to various mid-sized or small image recognition benchmarks [24], which helped them achieve state-of-the-art results, KArAts were never pre-trained with such datasets to showcase their full potential. Due to limited computing resources, we could not perform large-scale training. We investigate the KArAts’ transfer capability by fine-tuning on smaller datasets like SVHN [54], Oxford Flowers 102 [59], and STL-10 [18] from their ImageNet-1K pre-trained weights; see dataset details in §B.1.1. Table 2 shows KArAts transfer well across all datasets and have comparable performance with vanilla ViTs, even when their ImageNet-1K performance was not equivalent.
(iv) Generalizability: Detection & Segmentation. To understand the robustness of the Fourier KArAt, we employ ViT backbones for detection and segmentation. We choose Mask RCNN [33]-based framework for this purpose and employ a feature pyramid network [47] based on ViT-Det [46]; it is non-trivial to implement Fourier KArAt, see §B.1.5. We trained ViT-Det [46] using ViT-Base backbone on the MS COCO [48] dataset for 50 epochs with two settings: (i) fine-tuning on the ImageNet-1K pre-trained weights on traditional softmax attention, and (ii) random initialization. Table 4 shows that for the fine-tuning on ImageNet pre-trained weights, Fourier-KArAt shows $\sim 4-13\%$ gap in average precision (AP) from its vanilla variant; their qualitative performance in Figure 9 is similar. Overall, for ViT-Det, Fourier-KArAt performs inferiorly to its conventional counterpart in detection and segmentation. Interestingly, Figure 10 shows, vanilla ViT-Det using softmax activation reaches its peak performance quickly (within 12 epochs), Fourier-KArAt delays in achieving it (within 30 and 45 epochs for $G_{1}B$ and $G_{1}U$, respectively), proving the incompatibility of softmax-based weights initialization. It is, however, still a good initialization for KArAt as the performance is better than random initialization.
Figure 4: 3D-visualization of Loss landscape for ViT-Tiny and ViT-Base along the two largest principal component directions of the successive change of model parameters. KArAt’s loss landscapes are significantly less smooth than traditional attention; spiky loss landscapes are undesirable in terms of optimization stability and generalizability of the resulting model. See Figure 11 for the loss contours and the optimizer trajectory.

### v3

@@ 4 Our Architecture
Different from learning activations for MLP, as in KAN and KAT, implementing the generic learnable attention incurs impractical computational and memory costs on present computing hardware and DL toolkits. For learnable activation, $\tilde{\sigma}:\mathbb{R}^{N}\to\mathbb{R}^{N}$, a full-rank operator, $\Phi^{i,j}\in\mathbb{R}^{N\times N}$} acts on each row ${\cal A}^{i,j}_{k,:}$ of ${\cal A}^{i,j}$, and produces $\tilde{\sigma}({\cal A}^{i,j}_{k,:}).$ These operations are extremely compute-heavy and have large memory footprints. E.g., ViT-Tiny dosovitskiy2020vit ; wu2022tinyvit ; steiner2022howtotrainvit has 5.53M parameters and requires nearly 0.9 GB of GPU memory to train. Implementing the learnable attention equation 1 in ViT-Tiny training with B-splines of order 5 and grid size 10 and Fourier basis with grid size 10 increases the parameter count to 30.68M and 39.06M, respectively. We could not train the version with B-splines for its humongous memory requirements, and B-spline computations are non-parallelizable. The one with the Fourier basis is parallelizable, but it takes approximately 60 GB of GPU memory when computing with a batch of one image. This computational bottleneck is agnostic of the basis functions.
@@
Figure 2: Different configurations to update $\widehat{\Phi}$:(a) Blockwise configuration, where $\Phi^{i,1}\neq\Phi^{i,2}\neq\cdots\neq\Phi^{i,L}$ for all i=1,2,,hi=1,2,...,h; (b) universal configuration, where $\Phi^{i,1}=\Phi^{i,2}=\cdots=\Phi^{i,L}=\Phi^{i}$ for all i=1,2,,h.i=1,2,...,h.
@@
How can we remedy this? Deep neural networks exhibit low-rank structure oja1982simplified ; jain2016streaming . Recently, kwon2024efficient showed that across various networks, the training largely occurs within a low-dimensional subspace. Therefore, we postulate that the attention matrices also have an underlying low-rank structure.
@@
To validate this, we perform spectral analysis of the attention matrices before and after the softmax activation; see Figures 5-5. The scree test cattell1966scree shows that for traditional attention matrices without softmax, there are 8 significant singular values; after softmax activation, this count increases to 16. The above observation verifies that with or without the nonlinear activation, the attention matrix $\mathcal{A}^{i,j}$ possesses an underlying low-rank structure. This motivates us to use a lower-dimensional operator, $^{\widehat{\Phi}}$ for the learnable attention calculation with $r=12$. In §C.4.2-C.4.3, we perform an ablation to find the best $r.$
@@
Instead of $\Phi^{i,j}\in\mathbb{R}^{N\times N}$}, we use an operator, $\widehat{\Phi}^{i,j}\in\mathbb{R}^{r\times N}$} such that $r\ll N$}, and the new learned activation is $r$-dimensional for $k\in[N]$.  This process significantly reduces the computational overhead. Next, we post-multiply another learnable weight matrix, $W^{i,j}\in\mathbb{R}^{N\times r}$ to project them back to their original dimension. For each $k\in[N]$}, this operation results in computing $\widehat{\sigma}({\cal A}^{i,j}_{k,:})=\left[W^{i,j}\widehat{\Phi}^{i,j}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]\right]^{\top}$} (Figure 1(b)(iv)); see other variants in §C.5. Our approximation is not unique or the best; it is one of the feasible solutions given the present computing interface.
@@
Fourier Kolmogorov-Arnold Attention. Although the default basis for KANs is B-Splines in liu2024kan , the number of MHSA parameters is much more than that of an MLP. Specifically, for $L$ encoder blocks, each with $h$ attention heads, the parameter complexity is $O(N^{2}Lh(G+k))$, for $\Phi\in\mathbb{R}^{N\times N}$ and $O(NrLh(G+k))$, for $\widehat{\Phi}\in\mathbb{R}^{r\times N}$, if the model uses B-Splines of degree $k$.
@@
Moreover, as yang2024kolmogorov mentioned, B-splines are localized functions, not standard CUDA functions. Although efficient CUDA implementations for cubic B-Splines exist ruijters2012gpu ; ruijters2008efficient , their overall implementation results in slower, sparse, non-scalable, and complicated GPU execution; also, see xu2024fourierkan ; KANsLimits . In §C.4.1, our extensive experiments with B-spline basis in KAN for image classification tasks verify their poor generalizability on medium-scale datasets (e.g., CIFAR-10 and CIFAR-100).
@@
carleson1967convergence Let $f$ be an $L^{p}$} periodic function for $p\in(1,\infty]$}, with Fourier coefficients $\hat{f}(n)$}. Then $\lim_{N\to\infty}\sum_{|n|\leq N}{\hat{f}}(n)e^{inx}=f(x)$}, for almost every $x$.
In particular, if a function is continuously differentiable, its Fourier series converges to it everywhere. In the past, xu2024fourierkan ; mehrabian2024implicit used the Fourier basis in KAN, dong2024fan proposed a Fourier analysis network (FAN) to capture periodic phenomena. Motivated by these, we use the Fourier basis to approximate the effect of the smooth softmax function. We employ the Fourier basis, $\{\sin{(\cdot)},\cos{(\cdot)}\}$ with gridsize $G$ to design learnable attention. Algorithm 1 can be used optionally. For $\widehat{\Phi}^{i,j}=[\widehat{\phi}^{i,j}_{pq}]\in\mathbb{R}^{r\times N}$}, we have $\widehat{\Phi}^{i,j}:\mathbb{R}^{N}\to\mathbb{R}^{r}$}. Hence, each row ${\cal A}^{i,j}_{k,:}$}, for $k\in[N]$}, transformed into $\widehat{\Phi}^{i,j}_{p}\left[({\cal A}^{i,j}_{k,:})^{\top}\right]=\sum_{q=1}^{N}\widehat{\phi}^{i,j}_{pq}({\cal A}^{i,j}_{k,q})$}, via the Fourier bases such that
@@
where $m$ refers to the $m^{\rm th}$} grid point on a uniform grid of size $G$. The weights, $\{[a_{pqm}],[b_{pqm}]\}_{m=1}^{G},W^{i,j}]$} are updated in the backpropagation. If we use a Fourier basis in $\widehat{\Phi}\in\mathbb{R}^{r\times N}$}, the parameter complexity of MHSA becomes O(2NrhGL).{O(2NrhGL)}. Regardless of the basis used, the linear projection takes $O(N^{2}rhL)$ FLOPs. For any Transformer, $L$ and $h$ are fixed. So, one may omit them from the complexity results. See FLOPs and memory requirement in Table 13.
@@
[…]
@@
[…]
@@
[…]
@@
Blockwise and Universal operator configuration. We consider two configurations for updating the operator $\widehat{\Phi}^{i,j}.$ (a) Blockwise: In this configuration, each encoder block learns the attention through $h$ distinct operators $\widehat{\Phi}$ for each of the $h$ heads, totaling $hL$ operators; see Figure 2 (a). Like the MHSA in vanilla ViTs, the blockwise configuration is designed to learn as many different data representations as possible. (b) Universal: This is motivated from the KAT yang2024kolmogorov . In KAT, the MLP head is replaced with different variations of a KAN head—KAN and Group-KAN. In Group-KAN, the KAN layer shares rational base functions and their coefficients among the edges. Inspired by this, in our update configuration, all $L$ encoder blocks share the same $h$ operators; $\widehat{\Phi}^{i,j}=\widehat{\Phi}^{i}$ for $j=1,2,\ldots,L$; see Figure 2-(b). Rather than learning attention through $hL$ operators, this configuration only uses $h$ operators. Here, we share all learnable units and their parameters across $L$ blocks in each head. We postulate that the blockwise mode with more operators captures more nuances from the data. In contrast, the universal mode is suitable for learning simpler decision boundaries from the data. We also note that these two modes can be used with equation 1 as shown in Figure 1(b). Finally, Algorithm 2 in the Appendix provides the pseudocode for applying Fourier-KARAT in the ViT encoder block; see more nuanced discussion about them in §C.4.6.
@@ 5 Empirical Study
Implementing KArAt. We provide the implementation details in §C.1; see §C.1.1 for baselines and datasets. We integrate 5 wavelet bases, rational basis, Fourier basis, and 3 different base activation functions in our learnable MHSA design; see §C.1.3. Fourier KArAt performs the best, and we report it in the main paper. It has 1 hyperparameter, the grid size, $G$, and 2 configurations, blockwise or universal. We vary grid sizes from the set $G=\{1,3,5,7,9,11\}.$} For each value, we perform universal and blockwise updates; see Figure 2. With this formalization, $G_{n}B$ and $G_{n}U$ denote Fourier-KArAt with grid size $n$ and weights updated in blockwise and universal mode, respectively. See the ablation study of these variants in §C.5. We dispense the $\ell_{1}$ projection, as projecting the learned attention vectors to a probability simplex degrades the performance; see §C.4.5.
5.1 Trained Model Quality, Transferability, and Generalizability
Image Classification. Table 2 presents the performance of the best-performing Fourier KArAt variants; see Figure 6 for the training loss and test accuracy curves. Fourier KArAt is easily optimized in the initial training phase, gaining at par or better loss value and accuracy than the vanilla ViTs. However, in the later training phase, except for the fewer parameter variants (ViT-Tiny+Fourier KArAt), the loss curve flattens faster; we infer that for most of the Fourier-KArAt scenarios. We also notice that larger models like ViT-Base are more susceptible to slight changes in the parameters, making the training process challenging to manage, which we try to investigate in the rest of the paper using several analytical studies. Overall, ViT-Tiny+Fourier KArAt variants outperform ViT-Tiny on CIFAR-10 and CIFAR-100 by $5.40\%$ and $7.40\%$, respectively, and on ImageNet-1K, it achieves a similar accuracy. ViT-Small and -Base models with Fourier KArAt variants can barely outperform the vanilla models, and the accuracy differences increase as we tackle larger base models.
Performance on other ViTs for Image Classification. We fused KArAt with modern ViT-based architectures, ConViT convit and Swin Transformer liu2021swin ; §C.1.4 provides their implementation details. Table 4 shows that Fourier KArAt outperforms ConViT on CIFAR-10, experiences a modest drop on ImageNet-1K, and shows lower performance for hierarchical models, such as Swin, which requires a careful investigation.
Transferability: Image Classification. Unlike ViTs, which were pre-trained on large datasets and then transferred to various mid-sized or small image recognition benchmarks dosovitskiy2020vit , which helped them achieve state-of-the-art results, KArAts were never pre-trained with such datasets to showcase their full potential. Due to limited computing resources, we could not perform large-scale training. We investigate the KArAts’ transfer capability by fine-tuning on smaller datasets like SVHN svhn , Oxford Flowers 102 flowers102 , and STL-10 stl10 from their ImageNet-1K pre-trained weights; see dataset details in §C.1.1. Table 2 shows KArAts transfer well across all datasets and have comparable performance with vanilla ViTs, even when their ImageNet-1K performance was not equivalent.
Generalizability: Detection & Segmentation. To understand the robustness of the Fourier KArAt, we employ ViT backbones for detection and segmentation; see results and discussion in §C.2.
Table 3: Performance comparison of Fourier-KArAt and softmax attention on various vision transformer architectures.
@@ 5.1 Trained Model Quality, Transferability, and Generalizability
Table 4: Attention Transfer from ViT+Fourier-KArAt to vanilla ViTs, following li2024surprising .
Model Attention CIFAR-10 CIFAR-100
Transfered From Acc.@1 Acc.@5 Acc.@1 Acc.@5
ViT-Tiny None 72.76 98.14 43.53 75.00
[…]
ViT-Base None 83.45 99.19 58.07 83.70
[…]